Select the best possible first step to solving the system by first eliminating the x variable
x – 12y = –2
3x + 4y = 18

Answer by mananth(16949)

x – 12y = –2
3x + 4y = 18

(x – 12y = –2)*-3
-3x+36y=6................(1)
3x + 4y = 18 .............(2)
add the equations
40y= 24
y= 24/40
y=3/5
plug y in x-12y=-2
x-12(3/5)=-2
x-36/5 =-2
x=36/5 -2
x=26/5